# Simulating nationwide coupled disease and fear spread in an agent-based model

### v1

@@ Abstract
Human behavior and disease spread evolve together over the course of any disease outbreak. We present a dynamic model of that relationship based on the coupled spread of a disease and fear of that disease, and implement this model in a simulator, EpiCast, that simulates disease spread between individuals in a realistic synthetic population. In our model, fear of the disease spreads through both in-person contact and broadcast media, driving people to adopt behaviors that reduce disease spread. In order to better understand the dynamics of our model, we create and compare a range of compartmental surrogate models to analyze the impact of including various disease states. We also compare how Epicast-simulated outbreaks unfold across a number of scenarios with different types of behavioral responses to fear and a range of different parameter values. We find that the addition of asymptomatic, exposed, and pre-symptomatic disease states can impact both the rate at which an outbreak progresses and its overall trajectory. In addition, the combination of non-local fear spread through broadcasters and strong preventative measures by afraid individuals generally produces multiple epidemic waves, which only occur for a narrow range of parameter values with purely local fear spread.
keywords:
Agent-based model,Coupled contagions,Infectious diseases,Disease propagation1 Introduction
@@
Looking back on the early stages of the COVID-19 pandemic reveals a dizzying array of responses across all levels of society. From state-mandated lockdowns and school closures to individual masking, testing, and vaccination, these responses shaped not just personal experiences of the pandemic, but observed outcomes at a societal level as well [1, 2, 3, 4].
@@
Unfortunately, accurately anticipating individual behavior has proved challenging for the similarly diverse array of models developed to simulate the disease’s spread. The most common method of accounting for behavioral changes has been to use exogenous data sources, such as cell phone mobility data or historical public health policies, to impose top-down behaviors [5]. However, exogenous approaches struggle to capture the two-way feedback between behavioral changes and disease spread, since they mostly treat behavior as an input parameter, able to be scheduled in advance [6]. Truly dynamic behaviors require incorporating a behavioral model into the heart of a simulation, able to respond flexibly to changes in the global or local simulation state.
@@
However, most such endogenous behavioral models operate within high-level compartmental models where a series of ordinary differential equations (ODEs) govern the changing proportion of the population in different disease states [7]. Such models can only answer population-level questions about disease spread because they smooth over the heterogeneity of real-world populations. In contrast, agent-based models (ABMs) capture this heterogeneity by directly modeling the disease state and behavior of each individual in the simulated population. This capability has enabled the use of ABMs for retrospective analysis of COVID-19 spread patterns [8], evaluation of public health interventions [9, 10, 11], and the study of differential COVID-19 outcomes across regions and demographic groups [12]. However, ABMs rarely include endogenous behaviors [7], and the few existing ABMs which do only simulate relatively small populations consisting of at most a million people [13].
@@ 2 Related work
Coupled contagion models, where an individual’s behaviors depends on their beliefs, which are in turn informed by a combination of their neighbor’s beliefs and information regarding the overall state of the pandemic, either of which can spread in a manner similar to the disease itself. Epstein et al. present such a model where two contagions – fear of the disease and the disease itself – spread simultaneously through a population [19], adding a third contagion – fear of vaccination – in a later version of the model [20].
@@ 2.1 Agent-based modeling of endogenous behavior
Despite the range of developed behavioral models, the vast majority are built on top of standard compartmental models, most of which use a simple feedback loop. In particular, there are only a handful of agent based models (ABMs) which use coupled contagion models [7]. Epstein et al. develop a relatively simple ABM where 1800 agents move around a 2D grid and can spread fear and disease to agents in adjacent cells, and can either flee, isolate or retain the same behavior while fearful. Note that while they demonstrate the emergence of waves in the ordinary differential equation (ODE) version of their model, they fail to do so for the ABM [19]. Rajabi et al. also use a grid-based ABM, but only have agents quarantine when afraid. They demonstrate how multiple waves can emerge in their simulation and simulate the impact of travel restrictions and contact tracing on disease spread in a 5000 agent population [23]. Palomo-Briones et al. present a more complex but still spatially explicit 400 agent model where agents have a short visit schedule consisting of visits to their home, social activities, and economic activities, and can chose to wear masks, social distance, or wash hands based on a combination of other agents’ beliefs and the results of their previous behavior [24]. Curiel et al. use a small 5000-agent network-based ABM where a disease spreads alongside anti-vaccine beliefs under different network topologies and vaccination strategies [25]. Mao presents a detailed simulation of the Buffalo metropolitan area with nearly a million agents, with a schedule of activities drawn from demographic data. Agents can decide to use flu prophylaxis using a threshold-based decision model based on either infection risk or prophylaxis adoption among contacts [13]. While these models range in complexity, only Rajabi et al., with the simplest model, demonstrate an ability to produce multiple epidemic waves. In addition, the largest simulated population is smaller than most U.S. states and tuned for a narrow geographic area, which limits the usefulness of its results for state-level and national policy-makers.
@@ 2.2 The EpiCast simulator
In stark contrast to the smaller scale of the coupled contagion ABMS described above lie HPC-scale simulators such as EpiCast. EpiCast simulates disease spread on a massive artificial surrogate population of the full U.S. containing over 322 million agents, simulated people with individualized behavior [9, 14]. In the baseline version of EpiCast, this behavior relies on a combination of fixed exogenous responses based on demographics and top-down public health interventions taken as input. The population used in EpiCast is built using UrbanPop [26], a framework which uses American Community Survey (ACS) Data to create high-fidelity synthetic populations. Compared to similar frameworks, UrbanPop captures more of the heterogeneity in real-world populations, producing cross-sectional representations of agents across sociodemographic, economic, educational, and mobility characteristics. UrbanPop assigns each member of the generated population to a specific household in a specific census block group, in a manner which ensures the aggregate demographics of the generated block group correspond roughly to those observed in census data. In EpiCast, these households form the basis for dividing the population into communities, within which agents have varying contact intensities defined by whether they share a household, neighborhood, workplace, school, etc. These contact intensities determine the pairwise likelihood that disease transmission occurs between two agents within the same community. Agents then commute between a home and a work or school community each day in a manner consistent with Department of Transportation commute flow data, and occasionally engage in long distance travel to geographically distant communities based on flight data.
@@ 3 Methods
When designing our model, we set out to adapt an existing coupled contagion ODE model to function within EpiCast, which requires accounting for all of the disease states used in EpiCast. We first construct a SIR-based ODE-model as a simple baseline, then incrementally incorporate new disease states until we have accounted for all those present in the original EpiCast simulation. Once this design is complete, we then modify the model to function in the individualized context of EpiCast. In the process, we introduce both short- and medium-distance mechanisms for fear spread and two types of behavioral responses to fear with an eye toward reproducing the twin epidemic waves found in the ODE models.
@@ 3.1 Coupled contagion ODE models
We take a streamlined version of Epistein et al.’s model of coupled fear and disease spread citeepstein2008coupled as our starting point due to its flexibility and ability to produce multiple epidemic waves. We then incrementally add the asymptomatic, presymptomatic and exposed disease states present in EpiCast to this model. For brevity, we refer to this series of models using the Cartesian product of the two types of states. For example, we would refer to a standard model paired with a fear model with Neutral and Afraid states as an coupled model.
@@
, the closest to Epstein et al.’s original model. The flow rate from neutral and afraid compartments is slower by a factor of from the Recovered state. corresponds to the Epstein model.
@@
, distinguishing between symptomatic and asymptomatic Infectious and Recovered states. Of the disease states, only has a reduced flow rate from neutral to afraid compartments.
@@
For brevity, we present only the system in full. The flow diagram is shown in Fig S1 and the full system of equations is given in Section S1.
@@ 3.2.1 Modeling fear spread
We implement two methods of fear spread in EpiCast: one which operates based on an agent’s local neighborhood of contacts in a manner similar to disease spread, and a second that approximates the effect of traditional broadcast media.
In order to implement local fear spread, we took advantage of EpiCast’s preexisting methodology for computing disease spread, refactoring the code to allow for fear spread using different per-contact transmission probabilities. Fear is spread by symptomatic and afraid agents and countered by agents who know they have recovered from an infection and have a neutral fear state (i.e. agents in disease state and fear state ).
For non-local fear spread, we created broadcasters capable of taking one of three positions on the disease. They may either: (1) spread fear, (2) counter fear spread, or (3) remain neutral. In order to site these broadcasters, we begin by examining the occupations of agents in each community, which are identified by north American Industry Classification System (NAICS) code in EpiCast. If we identify at least one agent with a relevant NAICS code in a given community we locate a broadcaster there and assign all agents with the relevant occupations to work there. Each broadcaster then is able to potentially influence anyone within the census tract in which they are located, and starts by taking a neutral position.
Once initialized, a broadcaster’s position is determined by a combination of the fear states of its employees and the level of disease spread in the broadcast area. When the proportion of employees at a broadcaster who are afraid of the disease crosses a certain threshold, , the broadcaster begins spreading fear. We then record the current infection rate, the number of new infections in the broadcast area as a proportion of its population, as . In future timesteps, the broadcaster determines its position by comparing the current infection rate, , against as follows:
@@
Otherwise, the broadcaster spread fear of the disease.
@@ 3.2.2 Modeling behavior
We modeled two different ways in which agents may change their behavior when they are afraid: (1) taking protective measures (such as masking and hand-washing) while maintaining their normal routine, and (2) withdrawing from their normal routine and remaining at home. We model this first measure by reducing the susceptibility of agents in state , similar to how this is handled in [20] and in the ODE model.
In the second case, we modify an agent’s routine so that they remain in their home community, neither commuting to their work community nor engaging in long distance travel, and only interacting with other members of their household. If currently traveling, they immediately return to their home community. Note that this is different from a complete quarantine, as they can still transmit the disease or their fear to their household. If agents are afraid, they can either withdraw purely due to fear with a fixed probability for a given duration , or due to having a symptomatic infection (i.e. being in state ). Note that any agent with a symptomatic infection will withdraw if their symptoms are severe enough for them to be hospitalized, regardless of their fear state.
@@ 4.1 Experimental setup
First, we sought to compare solutions to each ODE model with an initial value problem (IVP) solver in order to identify scenarios under which each of our ODE models produce multiple epidemic waves, providing a comparison of when the same parameter values produce substantially different outcomes between models. This was intended to directly address RQ 1 and provide insights into which parameters might be of interest for RQ 3. Parameter values are shown in Table S1.
Second, we sought to compare several scenarios with different types of withdrawals – for hospitalized agents (hospitalized withdrawals), those who are symptomatic and afraid (sick withdrawals), and those who are merely afraid (fear-only withdrawals) – and both purely local (based on physical contact with agents in the same community) and non-local (based on broadcasters operating within a given census tract) fear spread modes in order to address RQ 2. For these scenarios, we use a synthetic population of about 322 million agents representing the 48 contiguous U.S. states and Washington DC. We seeded index cases on a per-county basis using estimates of active cases on 25 March 2020 derived from data from The New York Times, based on reports from state and local health agencies. For the first five days of the simulation, for an average of 856 thousand seed cases per day (about 0.265% of the population). Parameter values are shown in Table S2.
Finally, we investigated RQ 3 by conducting a series of sensitivity analyses comparing EpiCast outputs for a range of values of and , which determine the strength of an agent’s behavioral response to fear of the disease. We drew these values from the range , trying five values for each parameter. We ran each cell twice – once with only local fear spread and once with both local and broadcaster-based fear spread. These experiments were conducted on the contiguous U.S. dataset described above. Choices for parameters other than and correspond to Scenarios (c) and (d) from Table S2 (Or, equivalently, Scenarios (e) and (f)) without and with broadcasters, respectively. In order to compute the number of epidemic waves programmatically, we first compute a centered 7-day rolling maximum of new case counts – so as to avoid identifying peaks caused solely by weekly variation – then call the find_peaks function from the Scipy signal package (v1.14.0) in Python with a prominence of . This value indicates that new cases should fall by at least 1 new infection per 10,000 people before rising again if we are to consider a given local maximum to be the peak of a given epidemic wave. We found this value to avoid spurious peaks (i.e. those in the long tail of some outbreaks) while still capturing those most visually prominent in the data.
@@ 4.2 Comparing ODE models
For the ODE comparison experiments, Fig. 1 shows the solutions corresponding the parameter choices that either failed to produce a second epidemic wave within 360 days (top) or produced such a wave (bottom). For the model, we observe that changing the likelihood that those who recovered from a symptomatic infection become newly afraid of the disease, , has a dramatic impact on whether or not a second wave occurs. With the proportion of the population which are susceptible or afraid largely plateau around 0.8 and 0.6, respectively, after 360 days, as shown in Fig. 1(a). In contrast, these proportions plateau at around 0.35 and 0, respectively, when we use , as shown in Fig. 1(d). In additional, a second wave only occurs in the latter case, peaking a bit after 150 days.
When we introduce asymptomatic disease states, as in the model, the situation changes further. Keeping all parameters used in the run the same with this model results in 80% of the population still susceptible and 40% afraid after 360 days, and no second epidemic wave, as shown in Fig. 1(b). Changing the susceptibility of afraid agents, , from 25% to 35%, however, is sufficient to produce a second wave in this model, albeit a smaller one which peaks much later at around 275 days, as shown in Fig. 1(e). Under this scenario, about half of the population is still susceptible after 360 days although this proportion is still decreasing. The proportion afraid, while approaching zero by 360 days, does so about 160 days later than in the two wave scenario without asymptomatic infections (Fig. 1(d)).
Introducing an exposed state and a pre-symptomatic infectious state into the model produces the model, and further changes in the solution space. Using the same pair of parameter values as with the model results in faster disease spread and decay in fear levels. Fig. 1(c) shows that about 70% and 20% of the population is susceptible and afraid, respectively, after 360 days with in this model. Additionally, while we do not see a second wave peak within 360 days here, we do see the start of such a wave and the lack of a plateau of fear levels that corresponds to a single-wave outbreak in the other modes. When using with the model, Fig. 1(f) shows that fear levels reach zero and a second epidemic peak as with the model. However, both these events happen much sooner in the former model, although still about 100 days later than in the model.

### v2

@@ Abstract
Human cognitive responses, behavioral responses, and disease dynamics co-evolve over the course of any disease outbreak, and can result in complex feedbacks. We present a dynamic agent-based model that explicitly couples the spread of disease with the spread of fear surrounding the disease, implemented within the EpiCast simulation framework. EpiCast models transmission across a realistic synthetic population, capturing individual-level interactions. In our model, fear propagates through both in-person contact and broadcast media, prompting individuals to adopt protective behaviors that reduce disease spread. In order to better understand these coupled dynamics, we create and compare a range of compartmental surrogate models to analyze the impact of including various disease states. Additionally, we compare a range of behavioral scenarios within EpiCast, varying the level and intensity of fear and behavioral change. Our results show that the addition of asymptomatic, exposed, and pre-symptomatic disease states can impact both the rate at which an outbreak progresses and its overall trajectory. Moreover, the combination of non-local fear spread via broadcasters and strong behavioral responses by fearful individuals generally leads to multiple epidemic waves, an outcome that occurs only within a narrow parameter range when fear spreads purely through local contact. Accounting for the coupled spread of fear and disease is critical for understanding disease dynamics and designing timely, targeted responses to emerging infectious threats.
@@
Looking back on the early stages of the COVID-19 pandemic reveals a dizzying array of responses across all levels of society, across the entire world. The response to the COVID-19 pandemic may have been the largest coordinated behavior change in the history of humanity. From state-mandated lockdowns and school closures to individual masking, testing, and vaccination, these responses shaped both personal experiences of the pandemic, and also observed outcomes at a societal level [1, 2, 3, 4].
@@
Unfortunately, accurately anticipating individual behavior has proved challenging for the similarly diverse array of models developed to simulate the disease’s spread [5]. The most common method of accounting for behavioral changes has been to use exogenous data sources, such as cell phone mobility data or historical public health policies, to impose top-down behaviors [6]. However, exogenous approaches struggle to capture the two-way feedback between behavioral changes and disease spread, since they mostly treat behavior as an input parameter, able to be scheduled in advance [7]. Truly dynamic behaviors require incorporating a behavioral model into the heart of a simulation, able to respond flexibly to changes in the global or local simulation state.
@@
However, most such endogenous behavioral models operate within high-level compartmental models where a series of ordinary differential equations (ODEs) govern the changing proportion of the population in different disease states [8, 9]. Such models can only answer population-level questions about disease spread because they smooth over the heterogeneity of real-world populations. In contrast, agent-based models (ABMs) capture this heterogeneity by directly modeling the disease state and behavior of each individual in the simulated population. This capability has enabled the use of ABMs for retrospective analysis of COVID-19 spread patterns [10], evaluation of public health interventions [11, 12, 13], and the study of differential COVID-19 outcomes across regions and demographic groups [14]. However, ABMs rarely include endogenous behaviors [8], and the few existing ABMs which do only simulate relatively small populations consisting of at most a million people [15].
@@ 2 Related work
Coupled contagion models, where an individual’s behaviors depends on their beliefs. An individual’s beliefs are in turn informed by a combination of their neighbor’s beliefs and information regarding the overall state of the pandemic, either of which can spread in a manner similar to the disease. Epstein et al. present such a model where two contagions – fear of the disease and the disease itself – spread simultaneously through a population [21], adding a third contagion – fear of vaccination – in a later version of the model [22].
@@ 2.1 Agent-based modeling of endogenous behavior
Despite the range of developed behavioral models, the vast majority are built on top of standard compartmental models, most of which use a simple feedback loop. In particular, there are only a handful of agent based models (ABMs) which use coupled contagion models [8]. Epstein et al. [21] develop a relatively simple ABM where 1800 agents, simulated people with individualized behavior, move around a 2D grid. Agents can spread fear and disease to agents in adjacent cells, and can either flee, isolate or retain the same behavior while fearful. Note that while they demonstrate the emergence of waves in the ordinary differential equation (ODE) version of their model, they fail to do so for the ABM. Rajabi et al. [25] also use a grid-based ABM, but only have agents quarantine when fearful. They demonstrate how multiple waves can emerge in their simulation and simulate the impact of travel restrictions and contact tracing on disease spread in a 5000 agent population. Palomo-Briones et al. [26] present a more complex but still spatially explicit 400 agent model where agents have a short visit schedule consisting of visits to their home, social activities, and economic activities, and can chose to wear masks, social distance, or wash hands based on a combination of other agents’ beliefs and the results of their previous behavior. Curiel et al. [27] use a small 5000-agent network-based ABM where a disease spreads alongside anti-vaccine beliefs under different network topologies and vaccination strategies. Mao [15] presents a detailed simulation of the Buffalo metropolitan area with nearly a million agents, with a schedule of activities drawn from demographic data. Agents can decide to use flu prophylaxis using a threshold-based decision model based on either infection risk or prophylaxis adoption among contacts. While these models range in complexity, only Rajabi et al., with the simplest model, demonstrates an ability to produce multiple epidemic waves. In addition, the largest simulated population is smaller than most U.S. states and tuned for a narrow geographic area, which limits the usefulness of its results for state-level and national policy-makers.
@@ 2.2 The EpiCast simulator
The smaller scales of the coupled contagion ABMs described above contrast sharply with that of HPC-scale simulators like EpiCast. EpiCast simulates disease spread on a massive artificial surrogate population of the full U.S. containing over 322 million agents [11, 16]. In the baseline version of EpiCast, this behavior relies on a combination of fixed exogenous responses based on demographics and top-down public health interventions taken as input. The population used in EpiCast is built using UrbanPop [28], a framework which uses American Community Survey (ACS) Data to create high-fidelity synthetic populations. Compared to similar frameworks, UrbanPop captures more of the heterogeneity in real-world populations, producing cross-sectional representations of agents across sociodemographic, economic, educational, and mobility characteristics. UrbanPop assigns each member of the generated population to a specific household in a specific census block group, in a manner which ensures the aggregate demographics of the generated block group correspond precisely to those observed in census data. In EpiCast, these households form the basis for dividing the population into communities, within which agents have varying contact intensities defined by whether they share a household, neighborhood, workplace, school, etc. These contact intensities determine the pairwise likelihood that disease transmission occurs between two agents within the same community. Agents then commute between a home and a work or school community each day in a manner consistent with Department of Transportation commute flow data, and occasionally engage in long distance travel to geographically distant communities based on flight data.
@@ 3 Methods
In designing our model, we seek to adapt an existing coupled contagion ODE model to function within EpiCast, which requires accounting for all of the disease states present in the original code. We first construct a SIR-based ODE-model as a simple baseline, then incrementally incorporate new disease states until we have accounted for all those used by EpiCast. Once this design is complete, we then modify the model to function in an individualized context. In the process, we introduce both short- and medium-distance mechanisms for fear spread and two types of behavioral responses to fear with an eye toward reproducing the twin epidemic waves found in the ODE models.
@@ 3.1 Coupled contagion ODE models
We take a streamlined version of Epistein et al.’s model of coupled fear and disease spread [21] as our starting point due to its flexibility and ability to produce multiple epidemic waves. We then incrementally add the asymptomatic, presymptomatic and exposed disease states present in EpiCast to this model. For brevity, we refer to this series of models using the Cartesian product of the two types of states. For example, we would refer to a standard model paired with a fear model with Neutral and Fearful states as an coupled model.
@@
, the closest to Epstein et al.’s original model. The flow rate from neutral to fearful compartments is slower by a factor of from the Recovered state. corresponds to the Epstein model.
@@
, distinguishing between symptomatic and asymptomatic Infectious and Recovered states. Of the disease states, only has a reduced flow rate from neutral to fearful compartments.
@@
For brevity, we present only the system in full. The flow diagram is shown in Supplementary Fig. S1 and the full system of equations is given in Supplementary Section S1.
@@ 3.2.1 Modeling fear spread
[…]
In order to implement local fear spread, we took advantage of EpiCast’s preexisting methodology for computing disease spread, refactoring the code to allow for fear spread using different per-contact transmission probabilities. Fear is spread by symptomatic and fearful agents and countered by agents who know they have recovered from an infection and have a neutral fear state (i.e. agents in disease state and fear state ).
For non-local fear spread, we created broadcasters capable of taking one of three positions on the disease. They may either: (1) spread fear, (2) counter fear spread, or (3) remain neutral. In order to site these broadcasters, we begin by examining the occupations of agents in each community, which are identified by 2017 North American Industry Classification System (NAICS) codes in EpiCast. If we identify at least one agent who works in publishing (NAICS code 511) or broadcasting (NAICS code 515) in a given community we locate a media outlet there and assign all agents within the relevant industry to work there. Each broadcaster is able to potentially influence anyone within the census tract in which they are located, and starts by taking a neutral position.
Once initialized, a broadcaster’s position is determined by a combination of the fear states of its employees and the level of disease spread in the broadcast area. When the proportion of employees at a broadcaster who are fearful of the disease crosses a certain threshold, , the broadcaster begins spreading fear. We then record the current infection rate, the number of new infections in the broadcast area as a proportion of its population, as . In future timesteps, the broadcaster determines its position by comparing the current infection rate, , against as follows:
@@
Otherwise, the broadcaster spreads fear of the disease.
@@ 3.2.2 Modeling behavior
We modeled two different ways in which agents may change their behavior when they are fearful: (1) taking protective measures (such as masking and hand-washing) while maintaining their normal routine, and (2) withdrawing from their normal routine and remaining at home. We model this first measure by reducing the susceptibility of agents in state , similar to how this is handled by Epstein et al. [22] and in the ODE model.
In the second case, we modify an agent’s routine so that they remain in their home community, neither commuting to their work community nor engaging in long distance travel, and only interacting with other members of their household. If currently traveling, they immediately return to their home community. Note that this is different from a complete quarantine, as they can still transmit the disease or their fear to their household. If agents are fearful, they can either withdraw purely due to fear with a fixed probability for a given duration , or due to having a symptomatic infection (i.e. being in state ). Note that any agent with a symptomatic infection will withdraw if their symptoms are severe enough for them to be hospitalized, regardless of their fear state.
@@ 4.1 Experimental setup
First, we sought to compare solutions to each ODE model with an initial value problem (IVP) solver in order to identify scenarios under which each of our ODE models produce multiple epidemic waves, providing a comparison of when the same parameter values produce substantially different outcomes between models. This was intended to directly address RQ 1 and provide insights into which parameters might be of interest for RQ 3. Parameter values are shown in Supplementary Table S1.
Second, we sought to compare several scenarios with different types of withdrawals – for hospitalized agents (hospitalized withdrawals), those who are symptomatic and fearful (sick withdrawals), and those who are merely fearful (fear-only withdrawals) – and both purely local (based on physical contact with agents in the same community) and non-local (based on broadcasters operating within a given census tract) fear spread modes in order to address RQ 2. For these scenarios, we use a synthetic population of about 322 million agents representing the 48 contiguous U.S. states and Washington DC. We seeded index cases on a per-county basis using estimates of active cases on 25 March 2020 derived from data from The New York Times, based on reports from state and local health agencies. For the first five days of the simulation, for an average of 856 thousand seed cases per day (about 0.265% of the population). Parameter values are shown in Supplementary Table S2.
Finally, we investigated RQ 3 by conducting a series of sensitivity analyses comparing EpiCast outputs for a range of values of and , which determine the strength of an agent’s behavioral response to fear of the disease. We drew these values from the range , trying five values for each parameter for a total of 25 combinations. We ran each combination twice – once with only local fear spread and once with both local and broadcaster-based fear spread. These experiments were conducted on the contiguous U.S. dataset described above. Choices for parameters other than and correspond to Scenarios (c) and (d) from Supplementary Table S2 (or, equivalently, Scenarios (e) and (f)) without and with broadcasters, respectively. In order to compute the number of epidemic waves programmatically, we first compute a centered 7-day rolling maximum of new case counts – so as to avoid identifying peaks caused solely by weekly variation – then call the find_peaks function from the Scipy signal package (v1.14.0) in Python with a prominence of . This value indicates that new cases should fall by at least 1 new infection per 10,000 people before rising again if we are to consider a given local maximum to be the peak of a given epidemic wave. We found this value to avoid spurious peaks (i.e. those in the long tail of some outbreaks) while still capturing those most visually prominent in the data.
@@ 4.2 Comparing ODE models
For the ODE comparison experiments, Fig. 1 shows the solutions corresponding to the parameter choices that either failed to produce a second epidemic wave within 360 days (top) or produced such a wave (bottom). For the model, we observe that changing the likelihood that those who recovered from a symptomatic infection become newly fearful of the disease, , has a dramatic impact on whether or not a second wave occurs. With the proportion of the population which are susceptible or fearful largely plateau around 80% and 60%, respectively, after 360 days, as shown in Fig. 1(a). In contrast, these proportions plateau at around 35% and 0%, respectively, when we use , as shown in Fig. 1(d). In additional, a second wave only occurs in the latter case, peaking a bit after 150 days.
When we introduce asymptomatic disease states, as in the model, the situation changes further. Keeping all parameters used in the run the same with this model results in 80% of the population still susceptible and 40% fearful after 360 days, and no second epidemic wave, as shown in Fig. 1(b). Changing the susceptibility of fearful agents, , from 25% to 35%, however, is sufficient to produce a second wave in this model, albeit a smaller one which peaks much later at around 275 days, as shown in Fig. 1(e). Under this scenario, about half of the population is still susceptible after 360 days, although this proportion is still decreasing. The fearful proportion, while approaching zero by 360 days, does so about 160 days later than in the two wave scenario without asymptomatic infections (Fig. 1(d)).
Introducing an exposed state and a pre-symptomatic infectious state into the model produces the model, and further changes in the solution space. Using the same pair of parameter values as with the model results in faster disease spread and decay in fear levels. Fig. 1(c) shows that about 70% and 20% of the population is susceptible and fearful, respectively, after 360 days with in this model. Additionally, while we do not see a second wave peak within 360 days here, we do see the start of such a wave and the lack of a plateau of fear levels that corresponds to a single-wave outbreak in the other models. When using with the model, Fig. 1(f) shows that fear levels reach zero and a second epidemic peak occurs, as with the model. However, both these events happen much sooner in the former model, although still about 100 days later than in the model.
@@
Figure 1: ODE model comparison: We show outputs with attention to parameter choices which impact the emergence of multiple epidemic waves. On the top and bottom we show solutions corresponding to parameter choices which fail to produce, or produce, respectively, a second wave within 360 days. We show these results for the (1(a) and 1(d)), (1(b) and 1(e)), and (1(c) and 1(f)) models. For ease of comparison, we combine all infected (including ), recovered, and fearful compartments in each model into a single line in each of the plots above.